are only for non zero numbers. 10 Example Prove that -a xb= -(a x b) for a,bER. 10.1 Proof Proof
Answer #1

Here I'm using distributive law that is a x(b+c)=(axb)+(axc).also -a+a=0.solution is below thank you.

axb=-(axb) fer a, b Elr. prove that solna let consider a,b EIR Gaxb) + (axb) =(-ata) xb ... (by distributie - (low) =102xb ea